# Can Model Uncertainty Function as a Proxy for Multiple-Choice Question Item Difficulty?

### v1

@@ Abstract
Estimating the difficulty of multiple-choice questions would be great help for educators who must spend substantial time creating and piloting stimuli for their tests, and for learners who want to practice. Supervised approaches to difficulty estimation have yielded to date mixed results. In this contribution we leverage an aspect of generative large models which might be seen as a weakness when answering questions, namely their uncertainty, and exploit it towards exploring correlations between two different metrics of uncertainty, and the actual student response distribution. While we observe some present but weak correlations, we also discover that the models’ behaviour is different in the case of correct vs wrong answers, and that correlations differ substantially according to the different question types which are included in our fine-grained, previously unused dataset of 451 questions from a Biopsychology course. In discussing our findings, we also suggest potential avenues to further leverage model uncertainty as an additional proxy for item difficulty.
@@ 1 Introduction
In contrast to previous work, we explore a fundamentally different approach to MCQ item difficulty estimation. Specifically, we explore whether model uncertainty signals show correlation with student performance, with a special focus on choice order sensitivity, a model uncertainty metric unique to MCQs. Rather than testing LLMs’ capabilities using very large MCQ datasets as monolithic collections, we study students’ and LLMs’ behaviours at a finer-grained level, finding different degrees of correlations according to different aspects of the questions. Viewing LLMs as "simulated classroom environments" as we suggest in this work, paired with advancements in model interpretability metrics, contributes to devising novel approaches to the use of language technology for educational applications. [^1] All code will be made available upon acceptance.
@@ 3.1 Student Performance Dataset
For this study, a set of MCQs and student performance data, such as the proportion of students selecting each choice, is required. Unfortunately, such a dataset is not publicly available. Therefore, this study utilises data from the "Biopsychology" course at the Social Sciences Faculty of our own institution, covering content from the classic textbook "Biological Psychology" by Kalat (2016).
The dataset comprises $451$ MCQs. For each question, there are three possible answer choices, and the dataset records the fraction of students who selected each of those three options. The data was collected from eight examinations with an average of $268$ examinees ($SD=185$). Moreover, the questions were designed by the course instructor and have not been previously published, minimising the risk of data contamination.
@@ 4.1 Question-Level: Correlation of Entropy
Figure 3: Spearman Correlation between student and model choice entropy. Asterisks signify significant correlation, using a significance level of $\alpha=0.05$.
@@ 4.3 Choice-Level: Correlation of Rates
We further focus on the subset of correctly answered questions. Again, we do not observe consistent difference between the three choices. Additionally, in line with the previous analysis, the correlation for the questions of Type 2 significantly increases by only studying the subset of correctly answered questions, with the exception of the results of Mistral-7b. It is not immediately clear why Type 2 overall shows stronger correlation. We hypothesise that this is because "fill two blanks" questions are the most challenging for the tested models (cf. Table 2), so it is expected for the models to have lower certainty scores on those questions, which might be more likely to correlate with lower student selection rates.
@@ 5 Discussion and Conclusion
Shifting the focus to the bigger picture, through this research, we demonstrate that it can be beneficial to consider non-supervised learning approaches to tackle the task of MCQ item difficulty estimation. We especially focused on the evaluation of model certainty as a proxy for item difficulty, but alternative approaches should also be explored. As an example, future research can focus on the impersonation ability of LLMs, by evaluating whether system instructions (e.g., "Answer this question as if you were a bad student") can offer useful insights for difficulty estimation.
@@
In this study, we used a dataset of multiple-choice questions from the "Biopsychology" course at the Social Sciences Faculty of our institution. The data was aggregated across multiple students and anonymised, ensuring that individual student performance cannot be traced.

### v2

@@ Abstract
Estimating the difficulty of multiple-choice questions would be great help for educators who must spend substantial time creating and piloting stimuli for their tests, and for learners who want to practice. Supervised approaches to difficulty estimation have yielded to date mixed results. In this contribution we leverage an aspect of generative large models which might be seen as a weakness when answering questions, namely their uncertainty. Specifically, we exploit model uncertainty towards exploring correlations between two different metrics of uncertainty, and the actual student response distribution. While we observe some present but weak correlations, we also discover that the models’ behaviour is different in the case of correct vs wrong answers, and that correlations differ substantially according to the different question types which are included in our fine-grained, previously unused dataset of 451 questions from a Biopsychology course. In discussing our findings, we also suggest potential avenues to further leverage model uncertainty as an additional proxy for item difficulty.
@@ 1 Introduction
In contrast to previous work, we explore a fundamentally different approach to MCQ item difficulty estimation. Specifically, we explore whether model uncertainty signals show correlation with student performance, with a special focus on choice order sensitivity, a model uncertainty metric unique to MCQs. Rather than testing LLMs’ capabilities using very large MCQ datasets as monolithic collections, we study students’ and LLMs’ behaviours at a finer-grained level, finding different degrees of correlations according to different aspects of the questions. Viewing LLMs as "simulated classroom environments" as we suggest in this work, paired with advancements in model interpretability metrics, contributes to devising novel approaches to the use of language technology for educational applications. [^1] All code is publicly available:
https://github.com/LeonidasZotos/Uncertainty-as-Proxy-for-Difficulty.
@@ 2.1 Item Difficulty Estimation
Lastly, a study by Loginova et al. (2021) is perhaps most similar to our research. The authors explore using model certainty as a proxy for difficulty. However, their work involves presenting questions in pairs, with the task of selecting the more difficult one. Additionally, their focus is on language comprehension, which differs from our emphasis on factual knowledge assessment. Furthermore, while their focus is on Encoder-Only models, ours is on Decoder-Only models, which incorporate greater amounts of factual knowledge as a byproduct of their language modeling objective Zhao et al. (2023).
@@ 3.1 Student Performance Dataset
For this study, a set of MCQs and student performance data, such as the proportion of students selecting each choice, is required. Unfortunately, such a dataset is not publicly available. Therefore, this study utilises data from the "Biopsychology" course at the Social Sciences Faculty of the University of Groningen, covering content from the classic textbook "Biological Psychology" by Kalat (2016).
The dataset comprises $451$ MCQs. For each question, there are three possible answer choices, and the dataset records the fraction of students who selected each of those three options. The data was collected from eight examinations with an average of $268$ examinees (Standard Deviation of $185$). Moreover, the questions were designed by the course instructor and have not been previously published, minimising the risk of data contamination.
@@ 4.1 Question-Level: Correlation of Entropy
Figure 3: Spearman Correlation (implemented using the SciPy package Virtanen et al. (2020)) between student and model choice entropy. Asterisks signify significant correlation, using a significance level of $\alpha=0.05$.
@@ 4.3 Choice-Level: Correlation of Rates
We further focus on the subset of correctly answered questions. Again, we do not observe consistent difference between the three choices. Additionally, in line with the previous analysis, the correlation for the questions of Type 2 significantly increases by only studying the subset of correctly answered questions, with the exception of the results of Mistral-7b. It is not immediately clear why Type 2 overall shows stronger correlation. We hypothesise that this is because "fill two blanks" questions are the most challenging for the tested models (cf. Table 2), so it is expected for the models to have lower certainty scores on those questions, which might be more likely to correlate with lower student selection rates  [^11] As also mentioned earlier, the limited number of examples in this category does not allow for strong statements, but we observe this behaviour..
@@ 5 Discussion and Conclusion
Shifting the focus to the bigger picture, through this research, we demonstrate that it can be beneficial to consider non-supervised learning approaches to tackle the task of MCQ item difficulty estimation. We especially focused on the evaluation of model certainty as a proxy for item difficulty, but alternative approaches should also be explored. As an example, future research can focus on the impersonation ability of LLMs, by evaluating whether system instructions (e.g., "Answer this question as if you were a student that only studied half the book") can offer useful insights for difficulty estimation.
@@
In this study, we used a dataset of multiple-choice questions from the "Biopsychology" course at the Social Sciences Faculty of the University of Groningen. The data was aggregated across multiple students and anonymised, ensuring that individual student performance cannot be traced. Because of its format as provided to us, no further ethics approval was required from our institution to work with this dataset.
Acknowledgements
We would like to thank the Ubbo Emnius Foundation for the M20 Fund and the Jantina Tammes School for Digital Society, Technology and Artificial Intelligence for facilitating this research. We are also grateful for the contribution of Dr. Mark Nieuwenstein from the University of Groningen who provided us with the Biopsychology question set from his course. We are also thankful to the reviewers for taking the time to read and comment our paper.